# GraphPFN: A Prior-Data Fitted Graph Foundation Model

### v1

@@ 5.2 Experimental results
Table 2 and Table 3 present the results of our experiments. Below, we summarize and discuss our key observations. Additional experimental results are provided in Appendix B.
@@
The in-context learning performance of GraphPFN is promising. In particular, GraphPFN outperforms both classic GNNs and other GFMs on some datasets.
For example, on artnet-exp and city-reviews, GraphPFN (ICL) outperforms both baselines trained from scratch and G2T-FM applied in the ICL mode. Moreover, on artnet-exp, GraphPFN (ICL) even outperforms finetuned G2T-FM.
At the same time, performance on some datasets like hm-prices or city-roads-M is significantly worse than that of G2T-LimiX. We hypothesize that this is caused by a mismatch between our prior and the downstream datasets. For example, during pretraining, GraphPFN has seen only graphs with up to 194,425194{,}425 edges, while hm-prices contains 10,730,99510{,}730{,}995 edges. As another example, the city-roads-M dataset presents a traffic network, which has a significantly different topology from the graphs in our prior. However, we believe that poor performance of GraphPFN on these datasets is not a fundamental limitation of our approach, but rather a direction for future work. By extending the prior and making it more diverse, one can potentially improve the performance of GraphPFN on the datasets where it currently does not achieve the best results.
@@
On datasets with up to 50,00050{,}000 nodes, finetuned GraphPFN achieves state-of-the-art results.
In particular, on 7 out of 10 datasets where we were able to finetune GraphPFN, it achieved the best results across all considered methods, while having competitive performance on the remaining 3 datasets. Notably, on several datasets, such as artnet-exp, hm-prices, and artnet-views, finetuned GraphPFN outperforms the second-best method by more than two percentage points.
We hypothesize that such strong performance comes from the ability of GraphPFN to capture complex graph patterns via message passing, unlike G2T-FM, which is limited to hand-crafted graph-based features. However, we also note that the pretraining procedure plays a crucial role in this success. Replacing the pretrained graph adapters with randomly initialized ones and finetuning this model for a specific downstream dataset leads to significantly weaker results, as detailed in Appendix B.
@@ 6 Conclusion
In this work, we propose GraphPFN: a prior-data fitted graph foundation model for node-level tasks. Following the PFN framework, GraphPFN was pretrained on synthetic datasets drawn from our novel prior over attributed graphs to perform predictions in the in-context learning regime. Our experiments show promising results for GraphPFN. Even in the ICL regime, GraphPFN often outperforms classic GNNs and prior ICL GFMs. After finetuning, GraphPFN consistently achieves state-of-the-art results, sometimes bringing substantial improvements over the second-best method. We believe our work shows that pretraining graph foundation models on synthetic datasets drawn from a carefully designed prior can be a promising direction for developing truly generalizable graph foundation models. Despite promising results, the current implementation of GraphPFN has several limitations, which we discuss in Appendix A.
@@ Appendix A Limitations and future work
  • GraphPFN is difficult to scale to very large datasets since its current implementation requires processing the whole dataset at once. This leads to significant memory consumption, which prevents us from, for example, finetuning GraphPFN on some graphs from our benchmark. Developing more scalable graph foundation models can be a promising direction for future work. For example, one can utilize more memory-efficient TFMs (like TabICL (Qu et al., 2025)) as backbones or combine GraphPFN with sampling methods (Hamilton et al., 2017; Zeng et al., 2019; Zou et al., 2019; Chiang et al., 2019; Zeng et al., 2021; Balin & Çatalyürek, 2023).
@@
  • The proposed graph prior does not cover graphs from specific domains like traffic networks. We hypothesize that this can be a key reason for the degraded performance on some datasets. Extending the prior with more diverse graph random models (e.g., geometric graphs (Barthélemy, 2011)) can further improve the results of GraphPFN and make its performance more robust.
  • Due to substantial computational resources required to pretrain GraphPFN, our work has limited ablation. Investigating the importance of incorporating self-supervised objectives like masked graph modeling into the GraphPFN objective or analyzing different components of the graph prior can bring new insights to the field.
@@
  • GraphPFN inherits some limitations from its tabular backbone LimiX. For example, GraphPFN does not natively support handling more than 10 classes in multiclass classification. This limitation can be alleviated with further development of TFMs or by using approaches such as error-correcting output codes, as proposed in Hollmann et al. (2025).
@@
  • Currently, GraphPFN is limited to node-level prediction tasks and cannot handle link prediction or graph-level tasks (e.g., graph classification or regression). While we have taken a first step in this direction by adding a masked graph modeling head to GraphPFN during pretraining, GraphPFN still heavily relies on the presence of node-level labels to make predictions. Therefore, we cannot directly apply GraphPFN to link prediction tasks. Extending GraphPFN to other prediction tasks can be a promising direction for future research.
Appendix B Additional results
Since GraphPFN achieves strong performance in the finetuning regime, one may hypothesize that the performance comes solely from the powerful graph adapters, but not from the pretraining procedure. To test this hypothesis, we consider the following model. We start with LimiX and add graph adapters, so the architecture exactly matches that of GraphPFN. But instead of using pretrained weights for graph adapters, we employ random weights. In order to stabilize training, we initialized the last layers in all graph adapters with zeros, ensuring that the random initialization does not break the model. After that, we finetune this model following exactly the same protocol as GraphPFN. The results of this model and a comparison with GraphPFN are presented in Table 4. One can see that using random adapters instead of pretrained ones leads to significant drops in performance, supporting the importance of pretraining for achieving the strong performance of the finetuned GraphPFN.
Appendix C Synthetic graph examples
@@
We design our prior to generate graphs that are both realistic and diverse. In Figures 3 and 4, we provide examples of our synthetic graphs. Note that these graphs tend to exhibit both community structure and core-periphery structure. Our graph generation process allows us to easily control various graph properties such as their size or density.
Appendix D Implementation details
Since some datasets (specifically, amazon-ratings and questions) with text-embedding features have relatively high feature dimensionality, which prevented us from directly finetuning on these datasets, we applied PCA to reduce the feature dimensionality to 64. We did not apply PCA to the pubmed dataset, since in our preliminary experiments applying PCA to that dataset led to degraded performance.
@@ Appendix D Implementation details
tolokers-2 artnet-exp hm-prices city-roads-M artnet-views facebook amazon-ratings questions wiki-cs
GraphPFN (FT) $60.77\pm 0.72$ $53.08\pm 0.23$ $79.15\pm 0.33$ $65.65\pm 0.20$ $64.15\pm 0.21$ $93.06\pm 0.09$ $46.18\pm 0.17$ $20.50\pm 0.83$ $82.17\pm 0.22$
LimiX + GA (FT) $46.19\pm 0.73$ $46.79\pm 0.10$ $70.15\pm 0.44$ $63.22\pm 0.48$ $54.02\pm 0.57$ $90.82\pm 0.39$ $39.81\pm 0.40$ $15.88\pm 4.00$ $78.35\pm 0.73$
Figure 3: Example visualizations of denser graphs from our prior.
Figure 4: Example visualizations of sparser graphs from our prior.

### v2

@@ 5.2 Experimental results
Tables 1 and 2 present the results of our experiments. Below, we summarize and discuss our key observations. Additional experimental results are provided in Appendix C.
@@
GraphPFN shows strong ICL performance, outperforming GNNs and other GFMs on the GraphLand benchmark, while matching them on classic graph datasets.
Specifically, when compared with GNNs and other GFMs evaluated in an ICL regime, GraphPFN (ICL) achieves the best average rank on GraphLand datasets and only loses to GAT on classic datasets. Importantly, unlike G2T-FM, which also achieves strong ICL performance, GraphPFN does not require potentially heavy preprocessing like the computation of Laplacian positional encodings or PageRank.
@@
Ensembling further boosts the ICL performance of GraphPFN, in particular allowing it to outperform GNNs on average on both collections of datasets.
Specifically, GraphPFN (ICL, E) achieves better average rank than GNNs and G2T-LimiX (ICL, E) on both collections of datasets. When compared to TAG, GraphPFN outperforms it on all considered GraphLand datasets, and performs on par on classic datasets, while requiring no potentially heavy preprocessing.
Observation 3
Finetuning allows GraphPFN to achieve the best results on 10 out of 13 datasets, and the gains over the second-best method are often substantial.
Specifically, the finetuned and ensembled GraphPFN achieves the best performance among all considered methods on all datasets except tolokers-2, amazon-ratings, and pubmed. Moreover, GraphPFN often yields substantial gains over the second-best method, for example, of at least one percentage point on the artnet-exp, artnet-views, avazu-ctr, city-roads-M, hm-prices, twitch-views, and questions datasets.
The performance of finetuned GraphPFN without ensembling is also strong. Specifically, it achieves the second-best result across all considered methods on 7 out of 13 datasets, only losing to GraphPFN (FT, E). It also achieves better average ranks than all other methods except those evaluated in the (FT, E) regime on both collections of datasets.
We hypothesize that this strong performance stems from GraphPFN’s ability to capture complex graph patterns via message passing, in contrast to G2T-FM and TAG, which rely on hand-crafted graph-based features. At the same time, the pretraining procedure appears to be critical. When we replace the pretrained graph adapters with randomly initialized ones and finetune on a downstream dataset, performance drops substantially; see Appendix C for details.
Table 2: Evaluation results on classic graph datasets under a random $10\%/10\%/80\%$ train/val/test split. ICL stands for in-context learning, FT stands for fine-tuning, and E stands for ensembling. We report average precision for binary classification tasks and accuracy for multiclass classification tasks. We also report average rank (AR) in the last column. For each column, we highlight first, second, and third best results with a color.
amazon-ratings facebook pubmed questions wiki-cs AR
GCN $41.43\pm 0.49$ $91.26\pm 0.21$ $85.46\pm 0.19$ $15.42\pm 0.67$ $81.74\pm 0.21$ $\mathbf{11.4}$
GraphSAGE $40.07\pm 0.53$ $91.12\pm 0.22$ $86.04\pm 0.27$ $16.55\pm 0.64$ $81.50\pm 0.27$ $\mathbf{11.6}$
GAT $40.67\pm 0.55$ $92.61\pm 0.21$ $84.81\pm 0.24$ $16.75\pm 0.67$ $82.25\pm 0.27$ $\mathbf{9.2}$
GT $41.56\pm 0.40$ $91.71\pm 0.22$ $84.95\pm 0.19$ $14.03\pm 0.90$ $82.54\pm 0.21$ $\mathbf{10.4}$
OpenGraph (ICL) $29.36\pm 1.24$ $75.27\pm 5.05$ $70.30\pm 2.67$ $3.77\pm 0.65$ $75.66\pm 0.39$ $\mathbf{16.8}$
AnyGraph (ICL) $33.49\pm 3.44$ $61.17\pm 8.64$ $65.31\pm 6.26$ $4.27\pm 0.66$ $65.17\pm 2.51$ $\mathbf{17.0}$
TS-GNN (ICL) $43.00\pm 0.13$ $77.87\pm 2.73$ $64.41\pm 5.11$ $5.00\pm 0.48$ $46.25\pm 9.77$ $\mathbf{15.6}$
G2T-LimiX (ICL) $44.20\pm 0.26$ $91.30\pm 0.23$ $88.93\pm 0.46$ $15.32\pm 0.81$ $80.12\pm 0.23$ $\mathbf{10.2}$
GraphPFN (ICL) $42.09\pm 2.06$ $90.44\pm 0.37$ $89.79\pm 0.35$ $18.27\pm 2.24$ $79.58\pm 0.40$ $\mathbf{10.0}$
G2T-LimiX (ICL, E) $44.65\pm 0.18$ $91.66\pm 0.05$ $89.19\pm 0.18$ $15.46\pm 0.93$ $1.09\pm 0.09 8.8\mathbf${8.8}
TAG-TabPFNv2 (ICL, E) $43.97\pm 0.52$ $93.11\pm 0.17$ $87.80\pm 0.20$ $15.07\pm 1.32$ $82.59\pm 0.12$ $\mathbf{7.6}$
TAG-LimiX (ICL, E) $44.89\pm 0.15$ $92.85\pm 0.14$ $88.09\pm 0.18$ $15.53\pm 1.99$ $82.64\pm 0.17$ $\mathbf{5.8}$
GraphPFN (ICL, E) $44.52\pm 0.27$ $91.70\pm 0.21$ $90.36\pm 0.13$ $20.79\pm 0.49$ $81.69\pm 0.15$ $\mathbf{6.4}$
GCOPE (FT) $39.90\pm 0.43$ $85.08\pm 0.17$ $79.35\pm 0.70$ $6.59\pm 0.43$ $59.13\pm 1.20$ $\mathbf{15.6}$
G2T-LimiX (FT) $44.00\pm 0.92$ $92.17\pm 0.22$ $90.57\pm 0.13$ $20.16\pm 0.56$ $81.90\pm 0.43$ $\mathbf{6.0}$
GraphPFN (FT) $45.34\pm 0.35$ $92.91\pm 0.22$ $89.79\pm 0.35$ $21.85\pm 1.04$ $81.91\pm 0.44$ $\mathbf{4.0}$
G2T-LimiX (FT, E) $46.05\pm 0.23$ $92.53\pm 0.13$ $90.88\pm 0.14$ $20.83\pm 0.44$ $83.11\pm 0.39$ $\mathbf{2.6}$
GraphPFN (FT, E) $45.89\pm 0.53$ $93.94\pm 0.12$ $90.36\pm 0.13$ $22.76\pm 0.55$ $83.53\pm 0.25$ $\mathbf{1.6}$
@@ 6 Conclusion
In this work, we propose GraphPFN: a prior-data fitted graph foundation model for node-level tasks. Following the PFN framework, GraphPFN was pretrained on synthetic datasets drawn from a carefully designed prior over attributed graphs to perform predictions in the in-context learning setting. Our experiments show strong results for GraphPFN. Even in the ICL setting, GraphPFN often outperforms classic GNNs and prior ICL GFMs. After finetuning, GraphPFN consistently achieves state-of-the-art results, often bringing substantial improvements over the second-best method. We believe our work shows that pretraining graph foundation models on synthetic datasets drawn from a well-designed prior is a promising direction for developing truly generalizable graph foundation models. Despite promising results, the current implementation of GraphPFN has several limitations, which we discuss in Appendix A.
@@ Appendix A Limitations and future work
Despite promising results, the current implementation of GraphPFN has several limitations that may stimulate future research:
@@
  • GraphPFN is difficult to scale to very large datasets since its current implementation requires processing the entire dataset at once, which leads to significant memory consumption. Developing more scalable graph foundation models can be a promising direction for future work. For example, one can utilize more memory-efficient TFMs (like TabICL (Qu et al., 2025; 2026)) as backbones or combine GraphPFN with sampling methods (Hamilton et al., 2017; Zeng et al., 2020; Zou et al., 2019; Chiang et al., 2019; Zeng et al., 2021; Balin & Çatalyürek, 2023).
@@
  • The proposed graph prior is focused on social and information networks and does not cover graphs from specific domains such as traffic networks. Extending the prior with more diverse random graph models (e.g., geometric graphs (Barthélemy, 2011)) can further improve the results of GraphPFN and make its performance more robust.
@@
  • GraphPFN inherits some limitations from its tabular backbone LimiX. For example, GraphPFN does not natively support more than 10 classes in multiclass classification. This limitation can be alleviated with further development of TFMs or by using approaches such as error-correcting output codes, as proposed in Hollmann et al. (2025).
@@
  • Currently, GraphPFN is limited to node-level prediction tasks and cannot handle link prediction or graph-level tasks (e.g., graph classification or regression). While we have taken a first step towards link-level tasks by adding a masked graph modeling head to GraphPFN during pretraining, the model still heavily relies on the presence of node-level labels to make predictions, so we cannot directly apply GraphPFN to link prediction tasks.
Appendix B Graph prior analysis
We analyze the proposed graph prior from three complementary perspectives. First, we visually inspect the generated graph structures in Appendix B.1. Second, we provide a numerical comparison of graph characteristics to demonstrate the alignment between our synthetic prior and real-world datasets in Appendix B.2. Finally, we examine the effect of the specific graph generation procedure on downstream performance in Appendix B.3.
B.1 Synthetic graph examples
@@
We design our prior to generate graphs that are both realistic and diverse. In Figure 3, we provide examples of our synthetic graphs. Note that these graphs tend to exhibit both community structure and core-periphery structure. Our graph generation process allows us to easily control various graph properties such as graph size or density.
B.2 Numerical analysis of graph characteristics
To further assess how closely our synthetic datasets match the properties of real-world graphs, we perform a numerical analysis of several graph characteristics. Specifically, we sample 1,0001{,}000 synthetic datasets from the prior used in the first pretraining stage and compute the following statistics.
First, we report basic structural measures, including the average degree, the clustering coefficient, and an estimate of the average pairwise distance. Second, to quantify the tendency of adjacent nodes to have similar targets, we compute unbiased homophily (Mironov & Prokhorenkova, 2024) for classification datasets and assortativity (Newman, 2003) for regression datasets. Finally, to evaluate how well the graphs follow the power-law degree distribution, we compute the cumulative degree distribution: for each degree value $k$, we estimate $P_{k}=\mathbb{P}(\deg(u)\geq k)$. We then fit a linear regression to the pairs $(\log k,\log P_{k})$ and report the corresponding $R^{2}$.
Table 4 summarizes the distributions of these characteristics for synthetic datasets, while Table 3 reports the same statistics for real datasets. In most cases, the distribution of characteristics in the synthetic datasets covers that of the real-world datasets.
Table 3: Characteristics of real-world graphs from the GraphLand benchmark.
artnet-exp artnet-views avazu-ctr city-reviews city-roads-M hm-prices tolokers-2 twitch-views
Avg. degree $11.12$ $11.12$ $288.04$ $15.66$ $3.75$ $460.92$ $88.28$ $80.87$
Clustering coefficient $0.03$ $0.03$ $0.24$ $0.26$ $0.00$ $0.27$ $0.23$ $0.02$
Avg. pairwise distance $4.34$ $4.43$ $3.72$ $4.78$ $129.89$ $2.39$ $2.84$ $2.84$
Homophily $0.28$ $0.69$ $0.10$
Assortativity $0.19$ $0.18$ $0.74$ $0.12$ -0.41
Degree power-law $R^{2}$ $0.70$ $0.70$ $0.90$ $0.92$ $0.52$ $0.84$ $0.83$ $0.97$
Table 4: Summary of distribution of characteristics of synthetic graphs for the first pretraining stage.
Mean Std Min Q25 Median Q75 Max
Avg. degree $125.64$ $130.27$ $2.39$ $10.24$ $87.27$ $206.33$ $479.03$
Clustering coefficient $0.3$ $0.23$ $0.0$ $0.08$ $0.28$ $0.49$ $0.88$
Avg. pairwise distance $2.92$ $1.35$ $1.57$ $2.01$ $2.28$ $3.39$ $9.76$
Homophily $0.06$ $0.16$ 0.48-0.48 0.01-0.01 $0.0$ $0.08$ $0.96$
Assortativity $0.05$ $0.13$ 0.21-0.21 0.0-0.0 $0.0$ $0.07$ $0.7$
Degree power-law $R^{2}$ $0.68$ $0.21$ $0.11$ $0.51$ $0.7$ $0.86$ $0.99$
B.3 Pretraining on simpler random graphs
To assess the utility of the proposed graph generation procedure, we conduct an ablation study in which we replace it during pretraining with simpler random graph models. Specifically, we consider three alternatives: the Erdős–Rényi model (Erdős & Rényi, 1959), the degree-corrected SBM (Karrer & Newman, 2011), and the preferential attachment model (Albert & Barabási, 2002) with varying initial degrees. Due to the high computational cost, we run only the first stage of pretraining for all three alternatives. For a fair comparison, we evaluate them against an intermediate GraphPFN checkpoint taken after completing the first pretraining stage.
The results, presented in Table 5, indicate that our proposed procedure yields superior performance in most cases. The only exceptions are the Erdős–Rényi model on the city-roads-M dataset and the degree-corrected SBM on artnet-views. For artnet-views, the difference is small and can plausibly be attributed to stochasticity in the pretraining process. For city-roads-M, we hypothesize that the outcome reflects a mismatch between the structural biases of our procedure, which primarily produces graphs resembling information, social, and web networks, and road networks, which typically lack a pronounced community or core-periphery structure. In this setting, the Erdős–Rényi model, which does not enforce these properties, may be more suitable. Conversely, on most of the datasets, the results for the Erdős–Rényi model are worse than for the more complex approaches, which indicates that a complex and diverse prior is important for GraphPFN pretraining.
Table 5: Results of pretraining GraphPFN on simpler graph models.
artnet-exp artnet-views avazu-ctr city-reviews city-roads-M hm-prices tolokers-2 twitch-views
Erdős–Rényi $46.92$ $56.17$ $25.09$ $78.60$ $63.75$ $66.52$ $45.26$ $55.92$
Degree-corrected SBM $47.27$ $60.87$ $25.68$ $78.71$ $59.60$ $69.86$ $58.29$ $62.69$
Preferential attachment $48.16$ $58.56$ $28.09$ $76.88$ $59.14$ $69.16$ $55.53$ $61.78$
Ours $50.61$ $60.05$ $28.31$ $79.46$ $61.65$ $73.37$ $59.04$ $64.23$
Appendix C Additional results
@@
artnet-exp city-reviews tolokers-2 artnet-views avazu-ctr city-roads-M hm-prices twitch-views
GraphPFN (FT) $50.64\pm 1.27$ $80.49\pm 0.16$ $60.90\pm 0.98$ $63.90\pm 0.12$ $34.34\pm 0.50$ $66.04\pm 0.41$ $79.26\pm 0.43$ $78.17\pm 0.13$
Limix + GA (FT) $46.19\pm 0.21$ $78.94\pm 0.35$ $48.55\pm 1.77$ $53.84\pm 0.59$ $26.78\pm 0.16$ $63.41\pm 0.59$ $69.44\pm 0.48$ $74.01\pm 1.41$
Since GraphPFN achieves strong performance in the finetuning setting, one may hypothesize that the performance comes solely from the powerful graph adapters, but not from the pretraining procedure. To test this hypothesis, we consider the following model. We start with LimiX and add graph adapters, so the architecture exactly matches that of GraphPFN. But instead of using pretrained weights for graph adapters, we employ random weights. In order to stabilize training, we initialized the last layers in all graph adapters with zeros, ensuring that the random initialization does not break the model. After that, we finetune this model following exactly the same protocol as GraphPFN. The results of this model and a comparison with GraphPFN are presented in Table 6. One can see that using random adapters instead of pretrained ones leads to significant drops in performance, supporting the importance of pretraining for achieving the strong performance of the finetuned GraphPFN.
Appendix D Additional details on experimental setup
Data
Table 7 presents key statistics of datasets used in our main experiments. Importantly, the selected datasets cover diverse domains, spanning both homophilic and heterophilic structures, diverse feature types (both tabular and text-based) and tasks (both classification and regression).
Table 7: The key statistics of the considered graph datasets.
artnet-exp 50,40550{,}405 280,348280{,}348 $75$ $11.1$ cls. 2 no tabular
artnet-views 50,40550{,}405 280,348280{,}348 $50$ $11.1$ reg. no tabular
avazu-ctr 76,26976{,}269 10,984,07710{,}984{,}077 $260$ $288.0$ reg. no tabular
city-roads-M 57,07357{,}073 107,104107{,}104 $26$ $3.8$ reg. yes tabular
hm-prices 46,56346{,}563 10,730,99510{,}730{,}995 $41$ $460.9$ reg. no tabular
tolokers-2 11,75811{,}758 519,000519{,}000 $16$ $88.3$ cls. 2 no tabular
twitch-views 168,114168{,}114 6,797,5576{,}797{,}557 44 $80.9$ reg. no tabular
pubmed 19,71719{,}717 44,32444{,}324 $500$ $4.5$ cls. 33 yes text-based
questions 48,92148{,}921 153,540153{,}540 $301$ $6.3$ cls. 2 no text-based
wiki-cs 11,70111{,}701 215,603215{,}603 $300$ $36.9$ cls. 10 yes text-based
PCA
Since some datasets (specifically, amazon-ratings and questions) with text-embedding features have relatively high feature dimensionality, which prevented us from directly finetuning on these datasets, we applied PCA to reduce the feature dimensionality to 64. We did not apply PCA to the pubmed dataset, since, in our preliminary experiments, doing so led to degraded performance. As a result, GraphPFN ran out of memory during finetuning on pubmed, and we therefore report its in-context learning (ICL) performance for this dataset.
Ensembling
For G2T-FM and GraphPFN, we report test-time ensembling results. Specifically, we evaluate the same model under $10$ different random seeds and average the resulting predictions. The diversity among ensemble members is induced by several sources of stochasticity: (1) internal randomness of the TFM model (e.g., random positional encodings), (2) random shuffling of labels and input features, (3) PEARL in G2T-FM, and (4) random features in GraphPFN. In G2T-FM, we do not apply feature shuffling, because it degraded performance on some datasets in our preliminary experiments. When ensembling is combined with finetuning, we apply ensembling at every evaluation step, including validation evaluation for early stopping and the final test evaluation. However, during optimization we use a single randomly sampled ensemble member to compute predictions for each optimizer step.
Figure 3: Example visualizations of graphs from our prior.